FT MICROWAVE SPECTRUM OF SULFUR CHLORIDE FLUORIDE, ClSF

JÜRGEN PREUSSER and \underline{MICHAEL C. L. GERRY}, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver B. C., Canada, V6T 1Z1.

The microwave spectrum of sulfur chloride fluoride, ClSF, has been measured for the first time, using a pulsed-jet cavity Fourier-transform microwave spectrometer. The samples were prepared from mixtures of SCl_{2}/SF_{6} or Cl_{2}/SF_{6} in Ne carrier gas, using an electric discharge nozzle. Spectra of three isotopomers (^{35}Cl^{32}S^{19}F, ^{37}Cl^{32}S^{19}F, and ^{35}Cl^{34}S^{19}F) have been observed in natural abundance, and have been analysed to produce rotational constants, centrifugal distortion constants, and hyperfine parameters of Cl and F. The molecular geometry has been derived. The harmonic force field has been refined, and used as part of the structure determination. The Cl quadrupole coupling constants and F spin-rotation constants have given information on the electronic structure. In particular, some rather unusual, negative, F spin-rotation constants will be discussed in terms of a potential electronic spectrum.